# On the convergence of adaptive approximations for stochastic differential equations

### v4

@@ Appendix B Generating increments and integrals of Brownian motion
We emphasise that the results detailed in this section are already well-established in the SDE numerics literature [15, 36] and are only included here for completeness.
@@
Theorem B.3 (Conditional distribution of increments and space-time Lévy areas of Brownian motion within an interval, [15, Theorems 6.1.4 and 6.1.6]).
@@
Figure 4: Illustration of the Brownian increments and areas in Theorem B.3 (diagram from [15])
@@ Appendix C Norm of the determinant for a Gaussian random matrix
Let $p_{n}$ denote the probability density function of $|\det A_{n}\hskip 0.7113pt|$. Then the Mellin transform of $p_{n}$ is given by [10, Equation (2.3)], but with $a=\frac{1}{2}$ and $\beta=1$. That is,
@@
which is the $n$-th moment of a half-normal distribution (see [74, Equation (18)]).
@@
Although (70) immediately follows from [10, Equation (2.3)] and [74, Equation (18)], it was difficult for the authors to find it stated in the literature. Hence, (70) may be a lesser-known (but elegant) fact about the normal distribution.

### v5

@@ Appendix B Generating increments and integrals of Brownian motion
We emphasise that the results detailed in this section are already well-established in the SDE numerics literature [foster2020thesis, jelincic2024VBT] and are only included here for completeness.
@@
Theorem B.3 (Conditional distribution of increments and space-time Lévy areas of Brownian motion within an interval, [foster2020thesis, Theorems 6.1.4 and 6.1.6]).
@@
Figure 4: Illustration of the Brownian increments and areas in Theorem B.3 (diagram from [foster2020thesis])
@@ Appendix C Norm of the determinant for a Gaussian random matrix
Let $p_{n}$ denote the probability density function of $|\det A_{n}\hskip 0.7113pt|$. Then the Mellin transform of $p_{n}$ is given by [cicuta2000matrix, Equation (2.3)], but with $a=\frac{1}{2}$ and $\beta=1$. That is,
@@
which is the $n$-th moment of a half-normal distribution (see [winkelbauer2014moments, Equation (18)]).
@@
Although (70) immediately follows from [cicuta2000matrix, Equation (2.3)] and [winkelbauer2014moments, Equation (18)], it was difficult for the authors to find it stated in the literature. Hence, (70) may be a lesser-known (but elegant) fact about the normal distribution.